# New insight into the orbital parameters of the gamma-ray binary HESS J0632+057

### v1

@@ Abstract
gamma-rays: stars – binaries: spectroscopic – stars: emission-line, Be – stars: individual: HESS J0632+0571 Introduction
@@
Gamma-ray binaries are a rare sub-class of high-mass binaries where the majority of their non-thermal emission is produced at energies 1 MeV, in a distribution (see e.g. Dubus, 2013; Chernyakova et al., 2019). In all nine of the known gamma-ray binary systems, the companion star has been identified as either an O type star - LS 5039 (Clark et al., 2001), 1FGL J1018.6-5856 (Nolan et al., 2012), LMC P3 (Corbet et al., 2016), 4FGL J1405.1-611 (Corbet et al., 2019) and HESS J1832-093 (van Soelen et al., 2024) - or a Be type star - PSR J2032+4127 (Massey & Thompson, 1991), PSR B1259-63 (Johnston et al., 1994), LS I +61∘303 (Gregory & Taylor, 1978) and HESS J0632+057 (Høg et al., 2000) - with a compact object in the mass range of a neutron star or black hole. Two long-standing scenarios, involving either a pulsar-wind or microquasar, are routinely proposed to explain the very high energy (VHE) emission observed from these sources (Dubus, 2013). In the pulsar-wind scenario, the compact object is a young neutron star producing a relativistic pulsar wind, which collides with the stellar wind from the O/Be companion. This forms a stand-off shock, wherein particles are accelerated to very high energies. The non-thermal emission is then produced as the particles cool-down via synchrotron and inverse Compton processes in the shock. The microquasar scenario suggests that the VHE emission originates from particles accelerated to very high energies in a relativistic jet, which is produced via accretion onto the compact object. To date, the compact object has been identified as a young pulsar in three systems, namely, PSR B1259-63 (Johnston et al., 1992), PSR J2032+4127 (Lyne et al., 2015) and LS I +61∘303 (Weng et al., 2022). The possible detection of pulsed emission has also been reported for LS 5039 (Yoneda et al., 2020; Makishima et al., 2023) but has not been confirmed (Kargaltsev et al., 2023).
@@
The gamma-ray binary HESS J0632+057, situated near the Rosette Nebula, consists of a B0Vpe star (MWC 148) and an undetected compact object in a d orbit (derived from X-ray observations; Adams et al., 2021). Over the course of long-term X-ray and -ray monitoring, the X-ray and TeV light curves consistently displayed two maxima: the primary, sharper peak, occurs around phases (time of phase zero ; Falcone et al., 2010), while the secondary, flatter peak, occurs around phases (e.g. Adams et al., 2021). The interpretation of the multi-wavelength emission from HESS J0632+057 is complicated by the two different and incompatible orbital solutions reported by Casares et al. (2012) and Moritani et al. (2018) (hereafter referred to as C12 and M18, respectively). The C12 solution was derived from radial velocity (RV) measurements of the broad, weak absorption lines, phase-folded on the initial orbital period estimate of P d (Bongiorno et al., 2011). This orbital solution would place the phases coinciding with the two maxima in the light curves around apastron, which would be difficult to interpret. In the later study by M18, the RV measurements were instead obtained from the H emission line, and phase-folded on orbital periods of 308 and 313 d, derived from an independent periodicity search by the authors of optical and X-ray data. The M18 orbital solution would result in the primary X-ray/TeV peak occurring after apastron, while the secondary flatter peak occurs near periastron. They suggested that the maxima in the light curve are a result of the pulsar interacting with the denser equatorial circumstellar disc of the Be star, which is inclined with respect to the orbital plane, similar to what is proposed for PSR B1259-63 (see e.g. Chernyakova et al., 2015, 2021). However, given that the optical emission lines arise from the circumstellar disc and not directly from the photosphere of the B star, the line profile is affected by the structure and behaviour of the disc. Thus, any asymmetry or inhomogeneities in the disc could affect the RV measurements. In order to try to mitigate this, M18 measured their velocities from the H emission wings using the bisector method (Shafter et al., 1986). Since the wings arise from the disc’s inner region, nearest to the Be star surface, it is suggested that this region would be less affected by interactions with the pulsar, and would more closely follow the motion of the Be star.
@@
Given the very different orbital solutions found, the interpretation of the observed multi-wavelength emission has been discussed with reference to both the C12 and M18 solutions. Because of this uncertainty, in addition to the C12 and M18 orbital solutions, several other attempts to constrain the orbital parameters through emission modelling have been undertaken (see Malyshev et al., 2019; Tokayer et al., 2021; Chen & Takata, 2022; Kim et al., 2022). These orbital solutions are primarily based on assumptions concerning the origin of the observed X-ray/TeV flux peaks, either assuming the peaks are due to the pulsar-disc crossings (Malyshev et al., 2019; Chen & Takata, 2022) or that the primary and secondary peaks are as a result of the intra-binary shock (IBS) peaking at periastron, and Doppler beaming along the cometary tail, respectively (Tokayer et al., 2021; Kim et al., 2022).
@@
This paper is structured as follows: in Section 2 we present the optical spectroscopic observations obtained with SALT, the different methods implemented to measure the RVs, as well as the equivalent widths (EWs) and V/R ratios of the Balmer emission lines; in Section 3 we present the results from the SALT data and the implication of the new SALT data on the C12 and M18 orbital solutions, the behaviour of the circumstellar disc and the multi-wavelength emission. Finally the conclusions are presented in Section 4.
@@ 2.1 SALT observations and data reduction
Optical spectroscopic observations were undertaken using the High Resolution Spectrograph (HRS; Barnes et al., 2008) on SALT (Buckley et al., 2006) in HR mode (). The fibre-fed echelle spectrograph separates both the sky and object fibres into a blue-arm (3700-5550 Å) and red-arm (5550-8900 Å) which then undergo cross-dispersion onto the blue/red cameras.
Twenty-four (24) observations consisting of s (first semester) or s (subsequent semesters) exposures were obtained over the course of three observing semesters, between 2020 December 7 and 2023 March 11. The individual echelle orders of the sky and object fibres were extracted and wavelength calibrated by the SALT-HRS pipeline (Kniazev et al., 2016). The individual orders of the echelle spectra were then merged, with special care taken to minimize the noise where the different orders overlapped, as many of the broader spectral features extended across more than one echelle order. The merged sky fibre spectra were then subtracted from the corresponding merged object fibre spectra, before being normalised, and rebinned to ensure a uniform wavelength dispersion across the entire spectrum. Finally, the spectra were corrected to the barycentre using the standard procedures in iraf/pyraf, and a single nightly averaged spectrum was produced from the three exposures obtained each night, for both arms. Fig. 1 shows portions of the obtained blue- and red arm spectra, averaged over all observations. The spectrum of MWC 148 consists of strong Balmer emission lines - with either a fine structure at the line peak (H ) or a double peak structure (H H ) - and several low intensity Fe ii emission lines arising from the circumstellar disc of the Be star, in addition to the broad, weak absorption lines produced in the photosphere of the star.
@@ 2.2 Radial velocity measurements
In this work the RV measurements were obtained through two different methods. First, by fitting Voigt profiles to the wings of the H , H and H Balmer emission lines, and, second, through cross-correlation of regions of the spectra excluding the strong Balmer emission lines. Obtaining RV measurements from several different line species, corresponding to different regions across the circumstellar disc, should help mitigate any local variation within these different emission regions.
@@ 2.2.1 Radial velocity from the Balmer emission lines
We have made use of Voigt profile fitting to the wings of the H , H and H Balmer emission lines to obtain the RV measurements, similar to the method in the earlier study by Moritani et al. (2015). The wings are produced by the more central regions of the circumstellar disc, which should be less affected by any variability present in the disc from tidal interaction with the companion. We made use of the Voigt1D model in the astropy package (Astropy Collaboration et al., 2013a), in order to obtain a best-fit to the emission line data obtained with the standard least-squares procedure in scipy.
@@
(ii) Inspection of the spectra surrounding the emission lines showed a slight slope in the continuum around the H and H lines, while the H emission line emerges from the underlying absorption feature. Therefore, an additional continuum correction was performed on the selected wavelength region. For the H and H emission lines, a first order polynomial gives a good fit to the continuum. In the case of the H emission line, a fourth order polynomial was used to fit the continuum and the underlying absorption feature from the star.
(iii) Finally, the Voigt fitting was performed in a two-step process: First, an initial fit to each emission profile was performed, which excluded the central 5 Å region of the line, to avoid the double peak profile. Second, the Voigt profile was re-fit to only the wings of the emission line, which was defined as the data falling below 25 per cent of the peak of the initial fit.
@@ 2.2.2 Radial velocity from cross-correlation
We first attempted to measure the RVs from the broad absorption lines, by performing cross-correlation on spectral windows covering the individual He i (4121, 4387, 4471, 5047), Si iii (4567) and C iii/O ii/He i blend (4647-4651 Å) features. Crucially, this is similar to the method described in C12. The cross-correlation was performed using the xcsao package (Kurtz & Mink, 1998) in iraf/pyraf, against a template spectrum produced from the average of the available blue-arm observations, similar to the process used by e.g. Foellmi et al. (2003); Monageng et al. (2017); van Soelen et al. (2019). This allowed us to directly correlate the spectra with a template that matches the target and which is taken with the same instrumental configuration. However, this also introduces any instrumental effects into the template (e.g. the noise at the merger sites of the orders in echelle spectra) and causes the spectral features to be artificially broadened due to the RV off-set of each frame. In order to address the broadening, the template was created in a two step process: first the RV off-set between each spectrum and the average of all spectra was determined via cross-correlation; and second, the template was then produced by averaging over all spectra after the initial RV offset had been removed. The velocity of the template was determined from the weighted average of the RVs of the H and H emission lines, following the method discussed in Section 2.2.1. The RVs obtained for each spectral window via cross-correlation, as well as the average, are shown in the middle panel of Figure 2. Note, the errors shown are only the statistical errors obtained from the cross-correlation analysis. The RV curve obtained by C12 (re-folded on d) is shown by the dashed line. The new observations do not show an indication of the sharp minimum associated with their fit.
@@
Given that the attempts to cross-correlate predominantly against the broad absorption lines produced poor results, we focused on the narrower, lower intensity features in the spectra which are formed deeper in the photosphere and are less affected by the strong stellar wind, minimizing some of the scatter. These spectral regions also included narrow Fe lines which, while arising from the circumstellar disc, originate closer in than the Balmer lines (Arias et al., 2006; Zamanov et al., 2016a, b; Moritani et al., 2015). From the average peak separation of the emission lines in the SALT spectra we estimate the radii of the H , H , H and Fe ii emission regions in the disc to be , , and 10-13, respectively, where the stellar radius was taken as . The orbital separation at periastron, based on the existing C12 and M18 orbital solutions, ranges between 13.2-18 . Thus, at least, the Fe ii emission region, up to the H emission region, is within the separation distance and would be less affected by any tidal interaction with the compact object.
@@
The cross-correlation was performed separately on three regions of the blue spectrum, labelled R1 (4110-4290 Å), R2 (4460-4800 Å), and R3 (4910-5060 Å) in Fig. 1. These regions were chosen to exclude the Balmer emission lines and the broad interstellar line at Å. A template was produced following the process detailed above for each of the three cross-correlation regions. The xcsao package first determines the Fourier transform of the object spectrum and the conjugate of the Fourier transform of the template spectrum, before finding the cross-correlation from the product between the two (Kurtz & Mink, 1998). Before this is performed, a section of the high and low frequencies in the Fourier space are removed. This is typically undertaken so as to filter out noise (high frequency) and broader continuum features (low frequency) that are still present in the spectra (Kurtz & Mink, 1998). In our final cross-correlation, we increased the level of the low frequency filtering. This had the effect of minimizing the relative strength of the broad, weak absorption lines, while increasing the relative strength of the narrow features in the cross-correlation. This resulted in far less scatter in the RV measurements.
@@ 2.3 Variation of the emission lines
In addition to the RV measurements, we measured the variability in the EW and the V/R profile of the Balmer emission lines. The EW measurements were performed on H , H and H , by summing the area under the lines. The error in the EW was calculated following eq. 7 in Vollmann & Eversberg (2006), using a signal-to-noise ratio that was obtained using the method presented by Stoehr et al. (2008). The H and H EW values, when phase-folded, show evidence of an orbital modulation. There is also an indication of orbit-to-orbit variation, with the first semester observations being slightly larger than the second (see around phase 0.8). The trend in the phase-folded data shows a lower between , before a gap in the data not covered by the SALT observations, around the predicted phase of the periastron. The strength at is higher and steadily declines towards the ‘low’ state again.
@@ 3.2.1 Comparison with existing solutions
The average RV from our Voigt fitting (open red circles) and cross-correlation of regions R1, R2 and R3 (filled blue circles), are shown in the top panel of Fig. 4. The results are compared to the RV solutions obtained for the C12 and M18 data refolded on the newer orbital period of 317.3 d. The change in the orbital period does not introduce a significant change to the existing orbital solutions, as the orbital parameters are all within the uncertainties of the previously published solutions (see Table 2). However, we note that the eccentricity of the refolded M18 data increases to .

### v2

@@ Abstract
binaries: spectroscopic – stars: emission-line, Be – stars: individual: HESS J0632+057 – gamma-rays: stars.1 Introduction
@@
Gamma-ray binaries are a rare subclass of high-mass binaries, where the majority of their non-thermal emission is produced at energies 1 MeV in a distribution (see e.g. Dubus, 2013; Chernyakova et al., 2019). In all nine of the known gamma-ray binary systems, the companion star has been identified as either an O type star – LS 5039 (Clark et al., 2001), 1FGL J1018.6-5856 (Nolan et al., 2012), LMC P3 (Corbet et al., 2016), 4FGL J1405.1-611 (Corbet et al., 2019) and HESS J1832-093 (van Soelen et al., 2024) – or a Be-type star – PSR J2032+4127 (Massey & Thompson, 1991), PSR B1259-63 (Johnston et al., 1994), LS I +61∘303 (Gregory & Taylor, 1978) and HESS J0632+057 (Høg et al., 2000) – with a compact object in the mass range of a neutron star or black hole. Two long-standing scenarios, involving either a pulsar-wind or microquasar, are routinely proposed to explain the very high energy (VHE) emission observed from these sources (Dubus, 2013). In the pulsar-wind scenario, the compact object is a young neutron star producing a relativistic pulsar wind, which collides with the stellar wind from the O/Be companion. This forms a stand-off shock, wherein particles are accelerated to very high energies. The non-thermal emission is then produced as the particles cool-down via synchrotron and inverse Compton processes in the shock. The microquasar scenario suggests that the VHE emission originates from particles accelerated to very high energies in a relativistic jet, which is produced via accretion on to the compact object. To date, the compact object has been identified as a young pulsar in three systems, namely PSR B1259-63 (Johnston et al., 1992), PSR J2032+4127 (Lyne et al., 2015), and LS I +61∘303 (Weng et al., 2022). The possible detection of pulsed emission has also been reported for LS 5039 (Yoneda et al., 2020; Makishima et al., 2023) but has not been confirmed (Kargaltsev et al., 2023).
@@
The gamma-ray binary HESS J0632+057, situated near the Rosette Nebula, consists of a B0Vpe star (MWC 148) and an undetected compact object in a d orbit (derived from X-ray observations; Adams et al., 2021). Over the course of long-term X-ray and -ray monitoring, the X-ray and TeV light curves consistently displayed two maxima: the primary, sharper peak, occurs around phases – (time of phase zero ; Falcone et al., 2010), while the secondary, flatter peak, occurs around phases – (e.g. Adams et al., 2021). The interpretation of the multiwavelength emission from HESS J0632+057 is complicated by the two different and incompatible orbital solutions reported by (Casares et al., 2012, hereafter referred as C12) and (Moritani et al., 2018, hereafter referred as M18). The C12 solution was derived from radial velocity (RV) measurements of the broad, weak absorption lines, phase-folded on the initial orbital period estimate of d (Bongiorno et al., 2011). This orbital solution would place the phases coinciding with the two maxima in the light curves around apastron, which would be difficult to interpret. In the later study by M18, the RV measurements were instead obtained from the H emission line, and phase-folded on orbital periods of 308 and 313 d, derived from an independent periodicity search by the authors of optical and X-ray data. The M18 orbital solution would result in the primary X-ray/TeV peak occurring after apastron, while the secondary flatter peak occurs near periastron. They suggested that the maxima in the light curve are a result of the pulsar interacting with the denser equatorial circumstellar disc of the Be star, which is inclined with respect to the orbital plane, similar to what is proposed for PSR B1259-63 (see e.g. Chernyakova et al., 2015, 2021). However, given that the optical emission lines arise from the circumstellar disc and not directly from the photosphere of the B star, the line profile is affected by the structure and behaviour of the disc. Thus, any asymmetry or inhomogeneities in the disc could affect the RV measurements. In order to try to mitigate this, M18 measured their velocities from the H emission wings using the bisector method (Shafter et al., 1986). Since the wings arise from the disc’s inner region, nearest to the Be star surface, it is suggested that this region would be less affected by interactions with the pulsar, and would more closely follow the motion of the Be star.
@@
Given the very different orbital solutions found, the interpretation of the observed multiwavelength emission has been discussed with reference to both the C12 and M18 solutions. Because of this uncertainty, in addition to the C12 and M18 orbital solutions, several other attempts to constrain the orbital parameters through emission modelling have been undertaken (see Malyshev et al., 2019; Tokayer et al., 2021; Chen & Takata, 2022; Kim et al., 2022). These orbital solutions are primarily based on assumptions concerning the origin of the observed X-ray/TeV flux peaks, either assuming the peaks are due to the pulsar-disc crossings (Malyshev et al., 2019; Chen & Takata, 2022) or that the primary and secondary peaks are as a result of the intra-binary shock (IBS) peaking at periastron, and Doppler beaming along the cometary tail, respectively (Tokayer et al., 2021; Kim et al., 2022).
@@
This paper is structured as follows. In Section 2, we present the optical spectroscopic observations obtained with SALT, the different methods implemented to measure the RVs, as well as the equivalent widths (EWs) and V/R ratios of the Balmer emission lines; in Section 3, we present the results from the SALT data and the implication of the new SALT data on the C12 and M18 orbital solutions, the behaviour of the circumstellar disc and the multiwavelength emission. Finally, the conclusions are presented in Section 4.
@@ 2.1 SALT observations and data reduction
Optical spectroscopic observations were undertaken using the High Resolution Spectrograph (HRS; Barnes et al., 2008) on SALT (Buckley et al., 2006) in HR mode (). The fibre-fed echelle spectrograph separates both the sky and object fibres into a blue-arm (3700–5550 Å) and red-arm (5550–8900 Å) which then undergo cross-dispersion on to the blue/red cameras.
Twenty-four observations consisting of s (first semester) or s (subsequent semesters) exposures were obtained over the course of three observing semesters, between 2020 December 7 and 2023 March 11. The individual echelle orders of the sky and object fibres were extracted and wavelength calibrated by the SALT-HRS pipeline (Kniazev et al., 2016). The individual orders of the echelle spectra were then merged, with special care taken to minimize the noise where the different orders overlapped, as many of the broader spectral features extended across more than one echelle order. The merged sky fibre spectra were then subtracted from the corresponding merged object fibre spectra, before being normalized, and rebinned to ensure a uniform wavelength dispersion across the entire spectrum. Finally, the spectra were corrected to the barycentre using the standard procedures in iraf/pyraf, and a single nightly averaged spectrum was produced from the three exposures obtained each night, for both arms. Fig. 1 shows portions of the obtained blue- and red arm spectra, averaged over all observations. The spectrum of MWC 148 consists of strong Balmer emission lines – with either a fine structure at the line peak (H ) or a double peak structure (H H ) – and several low intensity Fe ii emission lines arising from the circumstellar disc of the Be star, in addition to the broad, weak absorption lines produced in the photosphere of the star.
@@ 2.2 Radial velocity measurements
In this work, the RV measurements were obtained through two different methods. First, by fitting Voigt profiles to the wings of the H , H and H Balmer emission lines, and, second, through cross-correlation of regions of the spectra excluding the strong Balmer emission lines. Obtaining RV measurements from several different line species, corresponding to different regions across the circumstellar disc, should help mitigate any local variation within these different emission regions.
@@ 2.2.1 Radial velocity from the Balmer emission lines
We have made use of Voigt profile fitting to the wings of the H , H and H Balmer emission lines to obtain the RV measurements, similar to the method in the earlier study by Moritani et al. (2015). The wings are produced by the more central regions of the circumstellar disc, which should be less affected by any variability present in the disc from tidal interaction with the companion. We made use of the Voigt1D model in the astropy package (Astropy Collaboration et al., 2013), in order to obtain a best-fitting to the emission-line data obtained with the standard least-squares procedure in scipy.
@@
(ii) Inspection of the spectra surrounding the emission lines showed a slight slope in the continuum around the H and H lines, while the H emission-line emerges from the underlying absorption feature. Therefore, an additional continuum correction was performed on the selected wavelength region. For the H and H emission lines, a first order polynomial gives a good fit to the continuum. In the case of the H emission line, a fourth order polynomial was used to fit the continuum and the underlying absorption feature from the star.
(iii) Finally, the Voigt fitting was performed in a two-step process: First, an initial fit to each emission profile was performed, which excluded the central 5 Å region of the line, to avoid the double peak profile. Secondly, the Voigt profile was re-fit to only the wings of the emission line, which was defined as the data falling below 25 per cent of the peak of the initial fit.
@@ 2.2.2 Radial velocity from cross-correlation
We first attempted to measure the RVs from the broad absorption lines, by performing cross-correlation on spectral windows covering the individual He i (4121, 4387, 4471, 5047), Si iii (4567), and C iii/O ii/He i blend (4647–4651 Å) features. Crucially, this is similar to the method described in C12. The cross-correlation was performed using the xcsao package (Kurtz & Mink, 1998) in iraf/pyraf, against a template spectrum produced from the average of the available blue-arm observations, similar to the process used by e.g. Foellmi et al. (2003), Monageng et al. (2017) and van Soelen et al. (2019). This allowed us to directly correlate the spectra with a template that matches the target and which is taken with the same instrumental configuration. However, this also introduces any instrumental effects into the template (e.g. the noise at the merger sites of the orders in echelle spectra) and causes the spectral features to be artificially broadened due to the RV off-set of each frame. In order to address the broadening, the template was created in a two step process: first the RV off-set between each spectrum and the average of all spectra was determined via cross-correlation; and second, the template was then produced by averaging over all spectra after the initial RV offset had been removed. The velocity of the template was determined from the weighted average of the RVs of the H and H emission lines, following the method discussed in Section 2.2.1. The RVs obtained for each spectral window via cross-correlation, as well as the average, are shown in the middle panel of Figure 2. Note, the errors shown are only the statistical errors obtained from the cross-correlation analysis. The RV curve obtained by C12 (re-folded on d) is shown by the dashed line. The new observations do not show an indication of the sharp minimum associated with their fit.
@@
Given that the attempts to cross-correlate predominantly against the broad absorption lines produced poor results, we focused on the narrower, lower intensity features in the spectra which are formed deeper in the photosphere and are less affected by the strong stellar wind, minimizing some of the scatter. These spectral regions also included narrow Fe lines which, while arising from the circumstellar disc, originate closer in than the Balmer lines (Arias et al., 2006; Zamanov et al., 2016a; Zamanov et al., 2016b; Moritani et al., 2015). From the average peak separation of the emission lines in the SALT spectra we estimate the radii of the H , H , H and Fe ii emission regions in the disc to be , , , and 10–13, respectively, where the stellar radius was taken as . The orbital separation at periastron, based on the existing C12 and M18 orbital solutions, ranges between 13.2 and 18 . Thus, at least, the Fe ii emission region, up to the H emission region, is within the separation distance and would be less affected by any tidal interaction with the compact object.
@@
The cross-correlation was performed separately on three regions of the blue spectrum, labelled R1 (4110–4290 Å), R2 (4460–4800 Å), and R3 (4910–5060 Å) in Fig. 1. These regions were chosen to exclude the Balmer emission lines and the broad interstellar line at Å. A template was produced following the process detailed above for each of the three cross-correlation regions. The xcsao package first determines the Fourier transform of the object spectrum and the conjugate of the Fourier transform of the template spectrum, before finding the cross-correlation from the product between the two (Kurtz & Mink, 1998). Before this is performed, a section of the high and low frequencies in the Fourier space are removed. This is typically undertaken so as to filter out noise (high frequency) and broader continuum features (low frequency) that are still present in the spectra (Kurtz & Mink, 1998). In our final cross-correlation, we increased the level of the low-frequency filtering. This had the effect of minimizing the relative strength of the broad, weak absorption lines, while increasing the relative strength of the narrow features in the cross-correlation. This resulted in far less scatter in the RV measurements.
@@ 2.3 Variation of the emission lines
In addition to the RV measurements, we measured the variability in the EW and the V/R profile of the Balmer emission lines. The EW measurements were performed on H , H , and H , by summing the area under the lines. The error in the EW was calculated following eq. 7 in Vollmann & Eversberg (2006), using a signal-to-noise ratio that was obtained using the method presented by Stoehr et al. (2008). The H and H EW values, when phase-folded, show evidence of an orbital modulation. There is also an indication of orbit-to-orbit variation, with the first semester observations being slightly larger than the second (see around phase 0.8). The trend in the phase-folded data shows a lower between , before a gap in the data not covered by the SALT observations, around the predicted phase of the periastron. The strength at is higher and steadily declines towards the ‘low’ state again.
@@ 3.2.1 Comparison with existing solutions
The average RV from our Voigt fitting (open red circles) and cross-correlation of regions R1, R2, and R3 (filled blue circles), are shown in the top panel of Fig. 4. The results are compared to the RV solutions obtained for the C12 and M18 data refolded on the newer orbital period of 317.3 d. The change in the orbital period does not introduce a significant change to the existing orbital solutions, as the orbital parameters are all within the uncertainties of the previously published solutions (see Table 2). However, we note that the eccentricity of the refolded M18 data increases to .